# OrbitZoo: Multi-Agent Reinforcement Learning Environment for Orbital Dynamics

### v1

@@ Appendix 0.D Experiments
For each experiment, we first provide a brief definition, highlighting key aspects of the mission. We then describe the environment setup, detailing the observation and action spaces, body characteristics, and reward functions. Finally, we present the results, assessing the generalization capabilities of each policy.
0.D.1 Learning Algorithm
@@
The chosen algorithm for all missions was PPO, since it is a state-of-the-art RL algorithm in environments with continuous action spaces, such as in the experiments carried out in this paper. The implementation was inspired by [8], which offers a minimal PPO setup for discrete and continuous action spaces. However, several modifications were introduced to enhance learning performance, as recommended in [3]. These include the use of generalized advantage estimation (GAE) instead of Monte Carlo estimation for calculating advantages, recalculating advantages at each epoch, and employing batch-based training. Other improvements include changes in the network architecture (with input normalization, tanh activation functions, and smaller initial weights), together with hyperparameter tuning and performance adjustments (such as calculation of expected returns only at training time).
@@
Layer Actor Network (Input: state_dim_actor) Critic Network (Input: state_dim_critic)
Input BatchNorm1d(state_dim_actor) BatchNorm1d(state_dim_critic)
@@
0.D.2 Hohmann Maneuver
@@
The Hohmann transfer experiment [22] was chosen due to its well-established relevance in orbital mechanics and its suitability as a benchmark for evaluating reinforcement learning frameworks in this domain. As one of the most fuel-efficient orbital transfer maneuvers, it provides a clear and analytically solvable problem that allows for direct comparison between RL-derived solutions and theoretical optima. Moreover, the Hohmann transfer incorporates key aspects of orbital dynamics, such as thrust application, trajectory optimization, and state transitions, making it an ideal testbed to validate the realism and effectiveness of OrbitZoo’s high-fidelity environment.
0.D.2.1 Definition
@@
Figure 12: Representation of the Hohmann Maneuver. A spacecraft is initially in an orbit with semi-major axis $R$. It applies a thrust $\Delta V$ which positions it on an eccentric orbit. When the spacecraft reaches apoapsis, it performs a second thrust $\Delta V^{\prime}$ which positions it on the intended red orbit, with a semi-major axis of $R^{\prime}$.
0.D.2.2 Environment Setup
We define a spacecraft with a mass of $m_{f}=200$ kg and $50$ kg of fuel, starting on a near-circular orbit with equinoctial elements $(a,e_{x},e_{y},h_{x},h_{y})=(2000+R_{E}\text{ km},0.007,0.006,0.041,0.015)$ and ending up on $(a,e_{x},e_{y},h_{x},h_{y})=(2030+R_{E}\text{ km},0.007,0.006,0.041,0.015)$, where $R_{E}=6378$ is the radius of Earth. Therefore, the objective is to raise the semi-major axis by 30 km.
According to these values, using the Hohmann transfer, the optimal maneuver requires $\Delta V=\Delta V^{\prime}=6.16$ m/s and a transfer time of $t_{H}=3826.1$ s. Therefore, we define that each episode contains 1000 steps of 5 seconds, giving extra time for the agents to reach the target. Instead of assuming instantaneous burns, which are unrealistic, the thrust is applied for the entire step (5 seconds). For this interval, the optimal forces are $F_{1}=308$ N and $F_{2}=307.9$ N in the along-track direction. The spacecraft is equipped with a thruster that provides a specific impulse of $I_{sp}=310$ s and is capable of performing thrusts in every direction.
For simplicity, we consider a Newtonian attraction model with no perturbations, and the agents to always start in the same anomaly with little uncertainty. Tthe agent completes its objective if it reaches the target orbit within a tolerance for every equinoctial element: (±$(\pm 100\text{ m},\pm 0.005,\pm 0.005,\pm 0.001,\pm 0.001)$.
@@ 0.D.2.2 Environment Setup
The reward function is designed to encourage the agent to apply large but accurate thrusts. First, the absolute difference between the current orbit at time $t$ and the target orbit, $\hat{s}$, is defined as $\Delta s_{t}=|s_{t}-\hat{s}|$. Next, the scalar progress is computed as $P_{t}=w^{T}\cdot(\Delta s_{t-1}-\Delta s_{t})/s_{t}$, where $w$ is a vector of constant weights assigned to each orbital element. The reward function is then formulated as:
@@
Initial Standard Deviation 0.4
@@
0.D.2.3 Results
The agent training hyperparameters are shown in Table 5. Additionally, two experiments were conducted to analyze the agent’s behavior, with two different reward functions as shown in Table 6. However, there was no extensive exploration to find optimal coefficients. For the second experiment, the generalization of the agent was tested by switching from a Newtonian attraction model to a spherical harmonic gravity, and using third body forces (Sun and Moon), solar radiation pressure and drag force.
@@ 0.D.2.3 Results
Experiment 1 $1$ 00
Experiment 2 $1$ $0.5$
The agent was trained until it successfully adopted a strategy to achieve the target orbit within the specified tolerance. Figure 13 illustrates that in both experiments, the agent achieved the target semi-major axis while also correcting some other orbital elements. However, in Experiment 1, the agent failed to bring the $h_{y}$ component within the tolerance. In contrast, the results of the second experiment show a closer approximation to the optimal maneuver. Notably, when using more realistic dynamics by adding perturbative forces, the agent is also able to reach the target orbit, even with a constant change in orbital parameters created by these perturbations.
@@
Overall, this experiment demonstrates that even when trained without perturbations, the policy successfully generalizes to real dynamics, adapting to complex gravity fields, third-body forces, solar radiation pressure, and atmospheric drag. This is particularly significant in LEO, where gravity and drag are the dominant influences on a satellite’s trajectory.
@@
0.D.3 Chase Target
0.D.3.1 Definition
@@ 0.D.3.2 Environment Setup
In this experiment, we have an agent/satellite, called follower, that starts on an orbit characterized by the following Keplerian elements $(a_{F},e_{F},i_{F},\omega_{F},\Omega_{F},M_{F})=(10000+R_{E}\text{ km},0.1.5.0^{\circ},10.0^{\circ}.10.0^{\circ}.10.0^{\circ})$. Additionally, there is a non-maneuverable body, called leader, that starts on an orbit with a much larger altitude and less inclination than the follower: $(a_{L},e_{L},i_{L},\omega_{L},\Omega_{L},M_{L})=(40000+R_{E}\text{ km},0.001.5.0^{\circ},10.0^{\circ}.10.0^{\circ}.10.0^{\circ})$. To avoid singularities, we use equinoctial parameters to represent the orbits of the follower $(a_{F},e_{x_{F}},e_{y_{F}},h_{x_{F}},h_{y_{F}})$ and the leader $(a_{L},e_{x_{L}},e_{y_{L}},h_{x_{L}},h_{y_{L}})$. The objective is for the follower to approximate the leader.
The satellite is characterized by being spherical with a radius of 5 m, a mass of 650 kg (from which 150 kg are fuel), and has a propulsion system containing a specific impulse of $I_{sp}=3000$ s. Finally, the dynamics for training are composed of Newtonian attraction without additional perturbations, and each episode contains steps of 70 seconds.
@@
The satellite thrust system is capable of performing thrusts in any direction with a maximum magnitude of 30 N, that is, $(T_{\text{max}},\theta_{\text{max}},\phi_{\text{max}})=(30,\pi,2\pi)$.
The reward function aims to minimize the error between each equinoctial element, including the anomaly. We first define the difference between each element $e$ as: $\Delta e=e_{L}-e_{F},\forall{e}\in\{a,e_{x},e_{y},h_{x},h_{y},M\}$. The final reward function weights each absolute difference by a factor $\alpha_{e}$:
$r_{t}=-\left(\frac{|\Delta a|}{a_{L}},|\Delta e_{x}|,\Delta e_{y}|,|\Delta h_{x}|,|\Delta h_{y}|,|\text{atan2}(\sin\Delta M,\cos\Delta M)|\right)\cdot\alpha^{T},$ (20)
where $\alpha=(\alpha_{a},\alpha_{e_{x}},\alpha_{e_{y}},\alpha_{h_{x}},\alpha_{h_{y}},\alpha_{M})$. The error for the semi-major axis is normalized by the target element (leader) due of its large scale, and the error in the anomaly needs to account for the periodical nature of the element, where errors vary between $[0,\pi]$.
0.D.3.3 Results
To assess the performance and generalization ability of the follower, we conducted three experiments incorporating all perturbative forces available in OrbitZoo, including harmonic gravity fields, third-body forces (Sun and Moon), solar radiation pressure, and atmospheric drag.

### v2

@@ Appendix E Experiments
For each experiment, we first provide a brief definition, highlighting the key aspects of the mission. We then describe the environment setup, detailing the objectives, environment characteristics, observation and action spaces, and reward functions. Finally, we present the results, assessing the generalization capabilities of each policy, followed by a discussion of the challenges encountered and potential improvements.
All experiments are fully supported by the accompanying code, which is available at the following repository: code repository.
E.1 Learning Algorithms and Architectures
E.1.1 PPO
@@
The chosen algorithm for a large part of the missions was PPO, since it is a state-of-the-art RL algorithm in environments with continuous action spaces, such as in the experiments carried out in this paper. The implementation was inspired by [8], which offers a minimal PPO setup for discrete and continuous action spaces. However, several modifications were introduced to enhance learning performance, as recommended in [4]. These include the use of generalized advantage estimation (GAE) instead of Monte Carlo estimation for calculating advantages, recalculating advantages at each epoch, and employing batch-based training. Other improvements include changes in the network architecture (with input normalization, tanh activation functions, and smaller initial weights), together with hyperparameter tuning and performance adjustments (such as calculation of expected returns only at training time).
@@
Layer Actor Network (Input: state_dim) Critic Network (Input: state_dim)
Input BatchNorm1d(state_dim) BatchNorm1d(state_dim)
@@
E.1.2 DDPG
DDPG was employed for the mission described in Section E.2, aiming to replicate the work presented in [40] within a custom environment. Our implementation builds upon [64], incorporating the Ornstein-Uhlenbeck noise process [76] for exploration, and enhanced with prioritized experience replay [62], adapted from the implementation in [63].
The architectures of both the actor and critic networks consist of two hidden layers with Tanh activation functions, as detailed in Table 5. No exhaustive hyperparameter tuning was performed. Unlike PPO, the DDPG critic network estimates the Q-value, evaluating the expected return of the specific action taken by the actor in a given state, hence receiving the state-action pair.
Table 5: Network Structure of DDPG Actor and Critic.
Layer Actor Network (Input: state_dim) Critic Network (Input: state_dim, action_dim)
Input BatchNorm1d(state_dim) BatchNorm1d(state_dim_critic)
Hidden Layer 1 Linear(state_dim, 512), Tanh Linear(state_dim_critic, 512 + action_dim), Tanh
Hidden Layer 2 Linear(512, 256), Tanh Linear(512 + action_dim , 256), Tanh
Output Linear(256, action_dim), Tanh Linear(256, 1)
E.1.3 DDQN
DDQN was employed as an alternative to PPO for the collision avoidance task (see Section E.5), following a similar approach to that in [13]. The implementation is based on the open-source project [18], which builds upon the original DQN algorithm introduced in [52]. In our implementation, we also incorporate prioritized experience replay [62] and a target Q-value network for increased stability.
The Q-network architecture consists of two hidden layers with Tanh activation functions, as detailed in Table 6. No extensive hyperparameter tuning was conducted to optimize performance.
Table 6: Network Structure of DDQN Q-value network.
Layer Q-Value Network (Input: state_dim)
Input BatchNorm1d(state_dim)
Hidden Layer 1 Linear(state_dim, 512), Tanh
Hidden Layer 2 Linear(512, 256), Tanh
Output Linear(256, action_dim)
E.2 OrbitZoo vs. SOTA: Environment Comparison
E.2.1 Definition
Many RL missions and environments developed by other authors (as shown in Table 1) are not publicly available, hindering reproducibility and making it challenging to assess the complexity of their implementation. Among the few exceptions is the work by D. Kolosa [40] and A. Herrera [31], whose code is openly accessible and supports the results presented. In this context, we replicate Kolosa’s first mission to demonstrate how OrbitZoo can be configured to achieve similar objectives—more easily and with greater flexibility. A reproduction of Herrera’s mission is included in the supplementary code.
E.2.2 Environment Setup
Objective and Environment Characteristics.
We design an environment inspired by Kolosa’s setup. The scenario involves a spacecraft with a dry mass of 500 kg and an initial fuel mass of 150 kg, equipped with a thruster of specific impulse $I_{\text{sp}}=3100$ s. The initial orbit is defined by the equinoctial elements $(a,e_{x},e_{y},h_{x},h_{y},M)=(5500+R_{E}~\text{km},0.153,0.128,0.041,0.015,10^{\circ})$, with the goal of reaching a target orbit characterized by $(a,e_{x},e_{y},h_{x},h_{y})=(6300+R_{E}~\text{km},0.154,0.171,0.042,0.019)$ – representing a general orbital transfer, where $R_{E}=6378$ is the radius of Earth. The dynamics follow Newtonian gravitational attraction. Each episode spans 692 steps, with a time interval of 500 s per step (approximately four days total).
Action Space.
Departing from Kolosa’s Cartesian thrust representation, we employ a polar thrust parameterization, as described in Section B.2. The action at each time step is defined as $a_{t}=(T,\theta,\phi)$, constrained by $(T_{\text{max}},\theta_{\text{max}},\phi_{\text{max}})=(0.6,\pi,2\pi)$.
Observation Space.
Building upon Kolosa’s observation space, we include an additional realistic and time-varying parameter: fuel mass, which decreases as thrust is applied. The observation at time $t$ is defined as $o_{t}=(s_{t},M_{t},f_{t})$, where $s_{t}=(a,e_{x},e_{y},h_{x},h_{y})$ are the equinoctial elements, $M_{t}$ is the mean anomaly, and $f_{t}$ is the remaining fuel.
Reward Function.
The reward function also follows the same formulation, computed based on the deviation between the current and target equinoctial elements. Specifically, for each element $e$, the difference $\Delta e$ is scaled by a corresponding weight $\alpha_{e}$, resulting in:
$r_{t}=-(\alpha_{a}\Delta a+\alpha_{e_{x}}\Delta{e_{x}}+\alpha_{e_{y}}\Delta e_{y}+\alpha{h_{x}}\Delta h_{x}+\alpha{h_{y}}\Delta h_{y}),$ (24)
where
$\Delta e=\begin{cases}\frac{\sqrt{(\hat{e}-e)^{2}}}{\hat{e}}&\text{if }e=a\\ \sqrt{(\hat{e}-e)^{2}}&\text{if }e\in\{e_{x},e_{y},h_{x},h_{y}\}\\ \end{cases}.$
E.2.3 Results
The weights used during training were kept identical to those in Kolosa:
$(\alpha_{a},\alpha_{e_{x}},\alpha_{e_{y}},\alpha_{h_{x}},\alpha_{h_{y}})=(10^{0},10^{0},10^{0},10^{1},10^{1}).$
The DDPG training hyperparameters are summarized in Table 7. The training performance using our DDPG implementation (see Section E.1.2) is presented in Figure 12.
Table 7: OrbitZoo vs. Kolosa: Training hyperparameters.
Parameter Value
Actor learning rate 0.00001
Critic learning rate 0.0001
Epochs 1
Discount factor ($\gamma$) 0.99
$\tau$ 0.01
$\mu$ 0
$\sigma$ 0.2
$\theta$ 0.15
$\Delta$ t 0.01
Memory Capacity 10 000
Initial Standard Deviation 0.5
Batch Size 256
Figure 12: OrbitZoo vs. Kolosa: Cumulative reward per episode using the implemented DDPG.
When comparing our training performance with Kolosa’s, we observe that our agent initially achieves a higher reward score but ultimately converges to a lower final score. This difference is likely due to architectural variations between the implementations. Nonetheless, the agent exhibits stable learning dynamics, with performance stagnating after approximately 200 episodes. Analyzing the evolution of the agent’s orbital elements (Figure 13), we find that the agent correctly learns the objective, as evidenced by the consistent reduction in error across all elements except $e_{y}$. However, similarly to Kolosa, the agent is unable to converge all orbital elements within the predefined tolerance and time constraints.
Figure 13: Orbital elements of the maneuver per step.
Moreover, OrbitZoo offers a distinctive feature: a real-time visualization interface that allows users to observe the agent’s maneuvers as they occur in space. This capability provides valuable insight into the learning dynamics and behavior of the agent, as illustrated in Figure 14.
This mission demonstrates that OrbitZoo can successfully reproduce similar RL missions with ease, while also offering additional configurations and features that facilitate the development and testing of RL agents.
(a) OrbitZoo environment.
(b) Kolosa’s environment. Extracted from [40].
Figure 14: Visual comparison of the maneuver in OrbitZoo and on Kolosa’s environment.
E.2.4 Challenges and Future Improvements
Since our DDPG implementation differs from the original, the agent learns a different strategy, resulting in diverging training outcomes. This highlights the need for further hyperparameter tuning to align the behavior more closely with expected results. Future work may focus on replicating additional environments and benchmark results, aiming to evaluate and enhance the generalization capabilities of OrbitZoo across a wider range of mission scenarios.
E.3 Hohmann Maneuver
@@
The Hohmann transfer experiment [33] was chosen due to its well-established relevance in orbital mechanics and its suitability as a benchmark for evaluating reinforcement learning frameworks in this domain. Furthermore, this maneuver is rarely addressed using reinforcement learning, highlighting the importance of investigating it. As one of the most fuel-efficient orbital transfer maneuvers, it provides a clear and analytically solvable problem that allows for direct comparison between RL-derived solutions and theoretical optima. Moreover, the Hohmann transfer incorporates key aspects of orbital dynamics, such as thrust application, trajectory optimization, and state transitions, making it an ideal testbed to validate the realism and effectiveness of OrbitZoo’s high-fidelity environment.
E.3.1 Definition
@@
Figure 15: Representation of the Hohmann Maneuver. A spacecraft is initially in an orbit with semi-major axis $R$. It applies a thrust $\Delta V$ which positions it on the eccentric orbit. Later, it performs a second thrust $\Delta V^{\prime}$ which positions it on the intended orbit, with a semi-major axis of $R^{\prime}$.
E.3.2 Environment Setup
Objective and Environment Characteristics.
We define a spacecraft with a mass of $m_{f}=200$ kg and $50$ kg of fuel, starting on a near-circular orbit with equinoctial elements $(a,e_{x},e_{y},h_{x},h_{y})=(2000+R_{E}\text{ km},0.007,0.006,0.041,0.015)$ and ending up on $(a,e_{x},e_{y},h_{x},h_{y})=(2030+R_{E}\text{ km},0.007,0.006,0.041,0.015)$. Therefore, the objective is to raise the semi-major axis by 30 km. The spacecraft is equipped with a thruster that provides a specific impulse of $I_{sp}=310$ s and is capable of performing thrusts in every direction.
According to the values provided, using the Hohmann transfer, the optimal maneuver requires $\Delta V=\Delta V^{\prime}=6.16$ m/s and a transfer time of $t_{H}=3826.1$ s. Therefore, we define that each episode contains 1000 steps of 5 seconds, giving extra time for the agents to reach the target. Instead of assuming instantaneous burns, which are unrealistic, the thrust is applied for the entire step (5 seconds). For this interval, the optimal forces are $F_{1}=308$ N and $F_{2}=307.9$ N in the along-track direction.
[…]
Action Space.
Observation Space.
@@ E.3.2 Environment Setup
Reward Function.
[…]
@@
Initial Standard Deviation 0.5
@@
E.3.3 Results
The agent training hyperparameters are shown in Table 8. Two experiments were conducted to analyze the agent’s behavior, with two different reward functions as shown in Table 9. However, there was no extensive research to find optimal coefficients.
@@ E.3.3 Results
Experiment 1 11 0
Experiment 2 11 $0.5$
For training, the weights used to measure the progress $P_{t}$ were $w=(10^{3},10^{0},10^{0},10^{1},10^{1},10^{-3})$, without extense research to find optimal parameters. The agent was trained until it successfully adopted a strategy to achieve the target orbit within the specified tolerance. Figure 16(a) shows the reward achieved on each experiment by episode, illustrating that Experiment 1 took more time to reach a good policy, explained by the lack of signaling regarding the along-track direction ($\alpha_{2}$). The fuel, as shown in Figure 16(b), shows that the agent learned to save fuel in both experiments.
(a) Cumulative Reward for Experiment 1 and Experiment 2.
(b) Final fuel for Experiment 1 and Experiment 2.
Figure 16: Hohmann Maneuver: Cumulative reward and fuel consumption across episodes.
For evaluation, in the second experiment, the generalization of the agent was tested by switching from a Newtonian attraction model to a spherical harmonic gravity, and using third body forces (Sun and Moon), SRP, and drag force.
Figure 17 illustrates that in both experiments, the agent achieved the target semi-major axis while also correcting some other orbital elements. However, in Experiment 1, the agent failed to bring the $h_{y}$ component within tolerance. In contrast, the results of the second experiment show a closer approximation to the optimal maneuver. Notably, when using more realistic dynamics by adding perturbative forces, the agent is also able to reach the target orbit, even with a constant change in orbital parameters created by these perturbations.
@@
Overall, this experiment demonstrates that even when trained without perturbations, the policy successfully generalizes to real dynamics, adapting to complex gravity fields, third-body forces, SRP, and atmospheric drag. This is particularly significant in LEO, where gravity and drag are the dominant influences on a satellite’s trajectory.
Additionally, the impact of providing a more informative reward function is evident, as Experiment 2 more quickly learns to avoid performing maneuvers outside the orbital plane.
E.3.4 Challenges and Future Improvements
Despite the agent’s ability to generalize to realistic dynamics, it still encounters difficulties reaching the target orbit consistently, as actions in PPO are stochastic by nature, whereas the Hohmann maneuver requires precise execution. Future improvements for this mission include employing reinforcement learning algorithms specifically designed for discrete actions, extending collision avoidance to multiple debris, and exploring hyperparameter tuning, since thrusts were not always applied at optimal moments.
E.4 Chase Target
E.4.1 Definition
@@ E.4.2 Environment Setup
Objective and Environment Characteristics.
In this experiment, we have an agent/satellite, called follower, that starts on an orbit characterized by the following Keplerian elements $(a_{F},e_{F},i_{F},\omega_{F},\Omega_{F},M_{F})=(10000+R_{E}\text{ km},0.1,5.0^{\circ},10.0^{\circ},10.0^{\circ},10.0^{\circ})$. Additionally, there is a non-maneuverable body, called target, that starts on an orbit with a much larger altitude and less inclination than the follower: $(a_{L},e_{L},i_{L},\omega_{L},\Omega_{L},M_{L})=(40000+R_{E}\text{ km},0.001,5.0^{\circ},10.0^{\circ}.10.0^{\circ},10.0^{\circ})$. To avoid singularities, we use equinoctial parameters to represent the orbits of the follower $(a_{F},e_{x_{F}},e_{y_{F}},h_{x_{F}},h_{y_{F}})$ and the leader $(a_{L},e_{x_{L}},e_{y_{L}},h_{x_{L}},h_{y_{L}})$. The objective is for the follower to approximate the leader.
The satellite is characterized by being spherical with a radius of 5 m, a mass of 650 kg (from which 150 kg are fuel), and has a propulsion system containing a specific impulse of $I_{sp}=3000$ s. Finally, the dynamics for training are composed of Newtonian attraction without additional perturbations, and each episode contains 2000 steps of 500 seconds.
Action Space.
The satellite thrust system is capable of performing thrusts in any direction with a maximum magnitude of 30 N, that is, limited by $(T_{\text{max}},\theta_{\text{max}},\phi_{\text{max}})=(30,\pi,2\pi)$.
Observation Space.
@@
Reward Function.
The reward function aims to minimize the error between each equinoctial element, including the anomaly. We first define the difference between each element $e$ as:
$\Delta e=\begin{cases}\frac{\sqrt{(e_{L}-e_{F})^{2}}}{e_{L}}&\text{if }e=a\\ \sqrt{(e_{L}-e_{F})^{2}}&\text{if }e\in\{e_{x},e_{y},h_{x},h_{y}\}\\ \text{atan2}(\sin{(e_{L}-e_{F}),\cos{(e_{L}-e_{F})}})&\text{if }e=M\end{cases}.$ (31)
The final reward function weights each difference by a factor $\alpha_{e}$:
$r_{t}=-\left(\alpha_{a}|\Delta a|+\alpha_{e_{x}}|\Delta e_{x}|+\alpha_{e_{y}}|\Delta e_{y}|+\alpha_{h_{x}}|\Delta h_{x}|+\alpha_{h_{y}}|\Delta h_{y}|+\alpha_{M}|\Delta M|\right).$ (32)
The error for the semi-major axis is normalized by the target element (leader) due of its large scale, and the error in the anomaly needs to account for the periodical nature of the element, where errors vary between $0,\pi]$.
E.4.3 Results
For training, the following weights were used, without extensive research for optimal parameters: $(\alpha_{a},\alpha_{e_{x}},\alpha_{e_{y}},\alpha_{h_{x}},\alpha_{h_{y}},\alpha_{M})=(10^{0},10^{-3},10^{-3},10^{-2},10^{-2},10^{-6})$. The evolution of training is shown in Figure 20.
Figure 20: Chase Target: Cumulative reward per episode.
To assess the performance and generalization ability of the follower, we conducted three experiments incorporating all perturbative forces available in OrbitZoo, including harmonic gravity fields, third-body forces (Sun and Moon), SRP, and atmospheric drag.